If I can spend ₹9 per month on electric light and the power is 30 paise per kWh. and I use 5 identical bulbs for 5 hours a day. What should be the power of each bulb?

Grade:10

2 Answers

yathartha gupta
5 years ago
Let 1 bulb =x kwh
 Hence,5 bulb=5x kWh
For 5 hours in 1 day = 5×5x = 25x kwh
 In a month =25x ×30 = 750x kwh
At the rate of 30 paisa per kwh the power will be
    750x × 0.30 = 225x rs
According to question ....
            225x = 9
         x = 9÷225
        x = 0.04 kW
     Thus,
                 The power of each bulb will be 0.04 kw
or 40 watt.
 
 
Ekta Agrawal
2 years ago
Let 1 bulb =x kwh
 Hence,5 bulb=5x kWh
For 5 hours in 1 day = 5×5x = 25x kwh
 In a month =25x ×30 = 750x kwh
At the rate of 30 paisa per kwh the power will be
    750x × 0.30 = 225x rs
According to question ....
            225x = 9
         x = 9÷225
        x = 0.04 kW
     Thus,
                 The power of each bulb will be 0.04 kw
or 40 watt.